OpenStudy (matt6288):

Evaluate -5x2 − 8x + 1 when x = -3. A. -20 # B. -10 C. 22 # D. 70

OpenStudy (anonymous):

c 22

jhonyy9 (jhonyy9):

A

jhonyy9 (jhonyy9):

-45+24+1=-45+25=-20